Circular No. 7780

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001io AND 2001ip G. Altavilla, S. Benetti, A. Pastorello, M. Riello, M. Turatto, and L.

Zampieri, Osservatorio Astronomico di Padova; E. Cappellaro, Osservatorio Astronomico di Capodimonte, A. Clocchiatti and C. Contreras, Universidad Catolica de Chile; I. J. Danziger and P. Mazzali, Osservatorio Astronomico di Trieste; and F. Patat, European Southern Observatory (ESO), report the discovery of two faint supernovae on CCD frames taken during a search performed with the ESO/MPI 2.2-m telescope (+ Wide Field Imager). The supernovae were identified by comparing the images with reference images taken (using the same equipment) on Nov. 11. Each object is within 1′′ of the nucleus of its host galaxy. Spectra obtained on Dec. 20 with the ESO 8.2-m MELIPAL Very Large Telescope (+ FORS1) reveal that both of these supernovae are of type Ia; from the spectral features, it is estimated that SN 2001io was 1–2 weeks after maximum and SN 2000ip was a few days before maximum.

SN  2001 UT  α_{2000}  δ_{2000}  V  z
2001io Dec. 8 0^{h}45^{m}53.^{s}64 −29^{o}26′11′′.7 23  0.19
2001ip Dec. 8 3 31 13.03 −27 50 55.5 23.8 0.54

SUPERNOVA 2001iq IN ANONYMOUS GALAXY G. M. Hurst, Basingstoke, England, reports the discovery by M. Arm-

strong, Rolvenden, England, of an apparent supernova (mag 17.2) on a 60-s unfiltered CCD image taken on Dec. 14.805 UT in the course of the U.K. Nova/Supernova Patrol with a 0.35-m Schmidt-Cassegrain reflector. The new object was at mag 16.6 on images obtained on Dec. 20.76, and it is located at α = 22^{h}25^{m}10^{s}.42, δ = +42^{o}43′04′′.2, (equinox 2000.0; average from two images), which is ≈ 3′′ southwest of a galaxy that itself is ∼ 87′′ west and 30′′ north of UGC 12032. Additional estimated magnitudes for SN 2001iq: 1989 Sept. 3, [20.8 (Palomar Observatory Sky Survey, red); Sept. 9, [22.5 (POSS, blue); 2001 July 22, [19.0 (Armstrong); Dec. 20.76, 16.6 (Armstrong).

COMET C/2001 W2 (BATTERS) Visual m_{1} estimates: Dec. 4.42 UT, 11.5 (K. Yoshimoto, Yamaguchi,

Japan, 0.25-m refl.); 9.72, 11.4 (W. Hasubick, Buchloe, Germany, 0.44-m refl.); 16.71, 11.8 (R. J. Bouma, Groningen, The Netherlands, 0.31-m refl.).

2001 December 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green